# What halts the growth of galaxies?

### v3

@@
The gas reservoir of galaxies can be altered by outflows driven by star-formation and luminous active galactic nuclei. Jets heating the surroundings of host galaxies can also prevent gas cooling and inflows. Spectacular examples for these three mass displacement channels have been observed, but their importance in transforming the galaxy population depends on the occurance rates of outflow triggers. We aim to investigate the absolute and relative importance of these three channels. In an observation-driven approach, we combine distribution functions and scaling relations to empirically compare average outflow rates across the galaxy total stellar mass spectrum and across cosmic time. This hinges on local outflow studies which should be extended to systematic, large and diverse samples, and we do not yet consider a halo heating effect by radiation-driven outflows. Our results show, independent of simulations, the dominance of star formation-driven outflows in low-mass galaxies. Massive galaxies today are predominately prevented from growing further by jet heating, while at $z=1-3$ all three processes are approximately similarly important. Over the full mass spectrum and cosmic history, outflows driven by the radiation from active galactic nuclei is never the dominant process.
@@ 2.3 Bubbles blown by active galactic nuclei
The resulting total mass rate is shown in the bottom right panel of Figure 1. We propagate the systematic uncertainties of the cavity-radio power relations, which vary most beween works, and the fit uncertainties in the mass-temperature and $M_{500}-M_{\star}$ relation. Unlike the outflows rates in the previous two sections, eq 2 computes the rate of an inflow that could have happened, were it not prevented by the AGN energy injection.
@@ 3 Results
4 Discussion & Conclusion
We presented a framework to compare key channels influencing the growth of galaxies. For this purpose, we combined state-of-the-art distribution functions and scaling relations of outflows driven by SF or AGN and cavities blown by AGN jets. Our approach is analoguous to semi-analytic modellers treatment of gas reservours that are filled by gravitational collapse. The gas reservoir can be emptied by outflows and gas accretion prevented by the heating of the galaxy inter-galactic medium (see e.g., Croton et al., 2006).
In our work, we have however not relied upon simulations. Instead, we obtained our results by combining observed relations collated from a large body of research undertaken over the last decades. As a prediction from Figure 3, we should see AGN-driven outflows dominate are wide-spread at $z\sim 1-2$, and at high galaxy stellar masses they dominate over SF-driven outflows. This is what is indeed found in high-resolution integral-field observations (e.g. Genzel et al., 2014).

### v4

@@
The gas reservoir of galaxies can be altered by outflows driven by star-formation and luminous active galactic nuclei. Jets heating the surroundings of host galaxies can also prevent gas cooling and inflows. Spectacular examples for these three mass displacement channels have been observed, but their importance in transforming the galaxy population depends on the occurrence rates of outflow triggers. We aim to investigate the absolute and relative importance of these three channels. In an observation-driven approach, we combine distribution functions and scaling relations to empirically compare average outflow rates across the galaxy total stellar mass spectrum and across cosmic time. This hinges on local outflow studies which should be extended to systematic, large and diverse samples, and we do not yet consider a halo heating effect by radiation-driven outflows. Our results show, independent of simulations, the dominance of star formation-driven outflows in low-mass galaxies. Massive galaxies today are predominately prevented from growing further by jet heating, while at $z=1-3$ all three processes are approximately similarly important. Over the full mass spectrum and cosmic history, outflows driven by the radiation from active galactic nuclei is never the dominant process.
@@ 2.3 Bubbles blown by active galactic nuclei
The resulting total mass rate is shown in the bottom right panel of Figure 1. We propagate the systematic uncertainties of the cavity-radio power relations, which vary most between works, and the fit uncertainties in the mass-temperature and $M_{500}-M_{\star}$ relation. Unlike the outflows rates in the previous two sections, eq 2 computes the rate of an inflow that could have happened, were it not prevented by the AGN energy injection.
@@ 3 Results
4 Discussion and conclusion
We presented a framework to compare key channels influencing the growth of galaxies. For this purpose, we combined state-of-the-art distribution functions and scaling relations of outflows driven by SF or AGN and cavities blown by AGN jets. Our approach is analogous to semi-analytic modelers treatment of gas reservoirs that are filled by gravitational collapse. The gas reservoir can be emptied by outflows and gas accretion prevented by the heating of the galaxy intergalactic medium (see e.g., Croton et al., 2006).
In our work, we have however not relied upon simulations. Instead, we obtained our results by combining observed relations collated from a large body of research undertaken over the last decades. As a prediction from Figure 3, we should see AGN-driven outflows dominate are widespread at $z\sim 1-2$, and at high galaxy stellar masses they dominate over SF-driven outflows. This is what is indeed found in high-resolution integral-field observations (e.g. Genzel et al., 2014).